OpenStudy (nottim):

How do you find the x-intercepts and y-intercepts of trinomials. E.g.(x^2-10x+25) How do you find the x-intercepts and y-intercepts of trinomials. E.g.(x^2-10x+25) @Mathematics

OpenStudy (anonymous):

well, for the y intercept simply set x to be 0 and solve for y. For the x intercpets i suggest you facrtor

OpenStudy (mertsj):

y=x^2-10x+25. find the x intercepts by setting y = to 0. That is: 0=x^2-10x+25. Solve that equation. That will be the x intercepts. If you get only 1 answer, that indicates a point of tangency to the x axis. If you get complex answers, that indicates that the parabola doesn't cross the x axis. Find the y intercept by setting x = 0. That is: y = (0)^2 -10(0) =25. That will be the y intercept.